# Steering Dialogue Dynamics for Robustness against Multi-turn Jailbreaking Attacks

### v1

@@ 4.3 Learning Invariant Safety Certificate to Defend Multi-turn Jailbreaking Attack
Since $\phi_{k}$ characterizes the evolving safe set ${\mathcal{X}}_{k}$ in the state space, it can reflect the satisfiability of the token-level responses $Z_{k}$ w.r.t the safety invariant subset of the user-specified region ${\mathcal{S}}_{I}\subset{\mathcal{S}}_{0}$ through the embedding mapping in Equation 2 and dialogue dynamics in Equation 3. Furthermore, the invariant safety condition in Equation 1 can be achieved through the following theorem, where the proof can be found in Section A.2.
@@
Given the neural dialogue dynamics in Equation 3 and the query embeddings $u_{k},k=1,2,\dots,K$, the LLM is invariantly safe according to Definition 3.1 if the following inequality conditions hold,
@@
In order to train the NBF with invariant safety conditions in Equation 8 to ensure LLM safety, the query context embedding set ${\mathcal{U}}_{k-1}$ needs to be quantified, which is challenging in the general multi-turn conversation cases. However, if the query context comes from multi-turn jailbreaking attack methods (Ren et al., 2024; Russinovich et al., 2024; Li et al., 2024b), we can assume the queries are adversarial — each query embedding $u_{k}$ results in the most unsafe predictions of $h$ given current context set ${\mathcal{U}}_{k-1}$ in both the current turn $k$ and the next turn $k+1$ — and then we have the following corollary to show the invariant safety conditions in Equation 8 empirically.
Corollary 4.3.
@@
then the invariant safety conditions in Equation 8 are satisfied if the following conditions hold,
@@
The proof of the corollary above can be found in Section A.3. We introduce the following empirical losses based on Equation 10: $\mathcal{L}_{SS}$ is the safe set loss to enforce the safe set satisfiability of $x_{k-1}\in{\mathcal{X}}_{k}$ based on the safety predictor, while $\mathcal{L}_{SI}$ is the safety invariance loss based on presumably invariant safety in the first $K-\kappa$ turns of dialogues, omitting non-invariantly-safe last $\kappa$ turns. Finally, combining Equations 4, 6, 11 and 12, the neural dialogue dynamics $f_{\theta},g_{\theta}$ and safety predictor $h$ can be jointly optimized as $\min_{f_{\theta},g_{\theta},h}\lambda_{dyn}{\mathcal{L}}_{dyn}+\lambda_{CE}{\mathcal{L}}_{CE}+\lambda_{SS}{\mathcal{L}}_{SS}+\lambda_{SI}{\mathcal{L}}_{SI}.$
@@
Based on the well-trained neural dialogue dynamics $f_{\theta},g_{\theta}$ and the NBF $\phi_{k}$ at each turn $k$, we introduce the filtering-based steering as a multi-turn jailbreak defense method. Given each state $x_{k-1}$ at each turn $k=1,2,\dots,K$, the NBF-based steering filters out all harmful attack queries $\hat{u}_{k}$ where $h(x_{k-1},\hat{u}_{k})+\eta\geq 0$ among jailbreaking context set ${\mathcal{U}}_{k-1}$ from attack methods, resulting in the safe query $u_{k}$ satisfying both Equation 9 and $\phi_{k}(x_{k-1})=h(x_{k-1},u_{k})<0$. Therefore, given the well-trained neural dialogue dynamics and NBF, with unsafe queries being filtered out by the proposed NBF-based steering, invariant safety condition in Equation 10 will be satisfied and LLM safety will not be jailbroken by Corollary 4.3. Empirically, the worst-case query candidates to be filtered come from jailbreaking attack methods (Ren et al., 2024; Russinovich et al., 2024; Li et al., 2024b), which are the most harmful reasonable queries along the context set ${\mathcal{S}}_{context}$ at each turn in Definition 3.1. An example of Crescendo attack (Russinovich et al., 2024) can be found in Figure 4, where $h(x_{k-1},u_{k})$ serves as a safety index to filter out the adversarial harmful query.
@@
LLMs LLMs w/ steering Attack Success Rate (ASR, $\downarrow$) Helpfulness ($\uparrow$)
ActorAttack Crescendo Opposite-day MTBench,1$\sim$10
GPT-3.5-turbo 0.585 0.560 0.785 8.00
GPT-3.5-turbo + steering 0.040 0.235 0.375 7.59
GPT-4o 0.600 0.565 0.725 9.35
GPT-4o + steering 0.035 0.260 0.325 8.77
o1 0.510 0.445 0.530 9.22
o1 + steering 0.090 0.280 0.210 8.83
Claude 3.5 Sonnet 0.200 0.215 0.095 9.14
Claude 3.5 Sonnet + steering 0.040 0.120 0.045 8.61
Table 1: Harmfulness and helpfulness of different closed-source LLMs.
@@ 5 Experiments
In this section, we aim to answer two questions: How does the proposed NBF-based safety steering perform as a defense method against different multi-turn LLM jailbreaking attacks on different LLMs? How is the proposed method influenced by steering threshold and training losses in terms of both safety and helpfulness? We answer the first question in Section 5.2 and the second one in Section 5.3, following the experimental setup. More details and results can be found in Appendix B.
@@ 5.1 Experimental Setup
To collect adversarial conversation data and safety labels for model training, we first generate diverse multi-turn jailbreaking attacks (Ren et al., 2024; Russinovich et al., 2024; Li et al., 2024b) and responses of GPT-3.5-turbo based on training queries from Circuit Breakers (Zou et al., 2024). Following Ren et al. (2024), we adopt GPT-4o as the LLM safety judge (Qi et al., 2023) to obtain safety scores ($1\sim 5$) as labels for each turn. Based on the collected multi-turn jailbreaking queries and responses, we first obtain the embeddings using the pre-trained Sentence-BERT model all-mpnet-base-v2 (Reimers and Gurevych, 2019) and use the embeddings to train the neural dialogue dynamics $f_{\theta},g_{\theta}$ based on Equation 4 for 200 epochs with Adam and learning rate $1e^{-4}$. We let the state dimension be 768, and $f_{\theta},g_{\theta}$ can be parameterized by 3-layer ReLU-based MLPs with the dimension shape of 1536-512-512-768. We then parameterize the safety predictor $h$ using 3-layer ReLU-based MLPs with dimension shape of 1536-32-32-5 based on the pre-trained neural dialogue dynamics $f_{\theta},g_{\theta}$. Given safety score labels, the predictor-based neural barrier function is learned based on Equations 6, 11 and 12 for 200 epochs with Adam and learning rate $1e^{-3}$. The steering threshold is $\eta=0$ and the number of non-invariant turns $\kappa$ is 3 by default during model training.
Safety and Helpfulness original
+ safe system
@@
llama-3-8b-instruct Attack Success Rate (ASR, $\downarrow$) ActorAttack 0.425 0.280 0.070 0.120 0.040
@@
Phi-4 Attack Success Rate (ASR, $\downarrow$) ActorAttack 0.405 0.370 0.100 0.080 0.015
@@
llama-3-8b-instruct Helpfulness ($\uparrow$) MMLU 66.00 65.66 63.34 64.52 46.65
MTBench 7.96 8.13 7.52 7.90 7.42
Phi-4 Helpfulness ($\uparrow$) MMLU 78.49 78.67 76.77 76.68 56.09
MTBench 8.23 8.59 8.06 8.18 7.76
Table 2: Safety and helpfulness comparison with baselines on different open-source LLMs.
@@
Evaluation and baselines.
We evaluate filtering-based safety steering on different LLMs, including GPT-3.5 (gpt-3.5-turbo-0125) (OpenAI, 2023), GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2024a), o1 (o1-2024-12-17) (OpenAI, 2024b), Claude-3.5 (claude-3-5-sonnet-20241022) (Anthropic, 2024), Llama-3-8b-instruct (Dubey et al., 2024), and Phi-4 (Abdin et al., 2024). We evaluate the defense performance against state-of-the-art multiturn jailbreaking attacks (ActorAttack (Ren et al., 2024), Crescendo (Russinovich et al., 2024), and Opposite-day (Li et al., 2024b)) on a Harmbench dataset (Mazeika et al., 2024; Ren et al., 2024), using Attack Success Rate (ASR) metric as the ratio of successful jailbreaks judged by GPT-4o (Qi et al., 2023). All temperatures are set to be 0.7. Besides, we evaluate the helpfulness using MMLU (Hendrycks et al., 2020) and MTBench (Zheng et al., 2023), where MMLU assesses a model’s general knowledge across various subjects and reports the percentage of accuracy while MTBench specifically evaluates an LLM’s ability to handle multi-turn conversations in dialogue scenarios with scores from 1 to 10. We implement two defense baselines in open-source LLMs: supervised fine-tuning (SFT) with LoRA (Hu et al., 2021; Zheng et al., 2024) and the safe system prompt from Llama-2-Chat (Llama-2-Chat, 2023; Touvron et al., 2023).
@@ 5.2 Result Comparison regarding Safety and Helpfulness on Multiple LLMs
Safety Comparison.
From Table 1, we can see that our proposed NBF-based steering with threshold $\eta=1e^{-3}$ can significantly reduce ASR against all multi-turn jailbreaking attacks on different LLMs, showing the effectiveness and strong generalizability of the proposed neural dialogue dynamics and barrier function to unseen LLMs. Compared to the defense baselines of safe system prompt and LoRA SFT in Table 2, our steering defense has the lowest ASR under the threshold $\eta=1e^{-3}$. Notably, LoRA SFT can even lead to a higher ASR against the Opposite-day attack due to the presence of benign data in the fine-tuning process, which is consistent with the findings of Qi et al. (2023); He et al. (2024); Qi et al. (2025).
Helpfulness Comparison.
We compare the helpfulness metrics in Tables 1 and 2 to show the trade-off caused by safety alignment. In Table 2, it can be seen that the safe system prompt can contribute to helpfulness while other defense methods tend to compromise helpfulness. Specifically, our steering with threshold $\eta=0$ is more helpful than LoRA SFT in both MMLU and MTBench. Although a stronger steering with threshold $\eta=1e^{-3}$ will cause a more significant decrease in MMLU due to filtering out unseen general knowledge without context, it still maintains satisfactory conversation ability indicated by the slight decrease in MTBench evaluation in Tables 1 and 2.
@@ 5.3 Ablation Study
Since the size of safe set ${\mathcal{X}}_{k}$ is controlled by the steering threshold $\eta>0$ in Equation 7, the larger $\eta$ is, the stronger safety steering will be. From Figure 5, we can see the Perato front induced by the steering threshold, showing the trade-off between helpfulness and safety. With the additional safe system prompt, LLMs can be safer and more helpful, but the attack success rate of jailbreaking is still high. Compared to LoRA SFT, the proposed safety steering has a better trade-off and is more flexible, being either more helpful given safety or safer given helpfulness.
@@
We compare the steering performance of safety and helpfulness based on NBF trained without either safe set loss ${\mathcal{L}}_{SS}$ of Equation 11 or safety invariance loss ${\mathcal{L}}_{SI}$ Equation 12 in Table 3 under safety steering threshold $\eta=0$. We can find that with respect to safety, ablating ${\mathcal{L}}_{SS}$ or ${\mathcal{L}}_{SI}$ will mostly increase ASR, showing that these two proposed losses are essential to train neural barrier functions. As a trade-off, adding ${\mathcal{L}}_{SS}$ or ${\mathcal{L}}_{SI}$ will slightly hurt single-turn MMLU helpfulness, while multi-turn helpfulness on MTBench will be better with safety invariance loss ${\mathcal{L}}_{SI}$.
LLMs ${\mathcal{L}}_{SS}$ ${\mathcal{L}}_{SI}$ Attack Success Rate (ASR, $\downarrow$) Helpfulness ($\uparrow$)
ActorAttack Crescendo Opposite-day MMLU MTBench
GPT-3.5-turbo, steering $\checkmark$ $\times$ 0.385 0.555 0.705 67.64 7.68
$\times$ $\checkmark$ 0.205 0.555 0.740 67.79 8.01
$\checkmark$ $\checkmark$ 0.135 0.430 0.655 66.24 7.93
Llama-3-8b- instruct, steering $\checkmark$ $\times$ 0.245 0.385 0.370 65.80 7.66
$\times$ $\checkmark$ 0.175 0.470 0.310 65.96 7.97
$\checkmark$ $\checkmark$ 0.120 0.360 0.310 64.52 7.90
Table 3: Effectiveness of the proposed safe set loss ${\mathcal{L}}_{SS}$ in Equation 11 and safe invariance loss ${\mathcal{L}}_{SI}$ in Equation 12.

### v2

@@ 4.3 Learning Invariant Safety Certificate to Defend Multi-turn Jailbreaking Attack
Since $\phi_{k}$ characterizes the evolving safe set ${\mathcal{X}}_{k}$ in the state space, it can reflect the satisfiability of the token-level responses $Z_{k}$ w.r.t the safety invariant subset of the user-specified region ${\mathcal{S}}_{I}\subset{\mathcal{S}}_{0}$ through the embedding mapping in Equation˜2 and dialogue dynamics in Equation˜3. Furthermore, the invariant safety condition in Equation˜1 can be achieved through the following theorem, where the proof can be found in Section˜A.2.
@@
Given the neural dialogue dynamics in Equation˜3 and the query embeddings $u_{k},k=1,2,\dots,K$, the LLM is invariantly safe according to Definition 3.1 if the following inequality conditions hold,
@@
In order to train the NBF with invariant safety conditions in Equation˜8 to ensure LLM safety, the query context embedding set ${\mathcal{U}}_{k-1}$ needs to be quantified, which is challenging in the general multi-turn conversation cases. However, if the query context comes from multi-turn jailbreaking attack methods (ren2024derail, ; russinovich2024great, ; li2024llm, ), we can assume the queries are adversarial — each query embedding $u_{k}$ results in the most unsafe predictions of $h$ given current context set ${\mathcal{U}}_{k-1}$ in both the current turn $k$ and the next turn $k+1$ — and then we have the following corollary to show the invariant safety conditions in Equation˜8 empirically.
Corollary 4.2.1.
@@
then the invariant safety conditions in Equation˜8 are satisfied if the following conditions hold,
@@
The proof of the corollary above can be found in Section˜A.3. We introduce the following empirical losses based on Equation˜10: $\mathcal{L}_{SS}$ is the safe set loss to enforce the safe set satisfiability of $x_{k-1}\in{\mathcal{X}}_{k}$ based on the safety predictor, while $\mathcal{L}_{SI}$ is the safety invariance loss based on presumably invariant safety in the first $K-\kappa$ turns of dialogues, omitting non-invariantly-safe last $\kappa$ turns. Finally, combining Equations˜4, 6, 11 and 12, the neural dialogue dynamics $f_{\theta},g_{\theta}$ and safety predictor $h$ can be jointly optimized as $\min_{f_{\theta},g_{\theta},h}\lambda_{dyn}{\mathcal{L}}_{dyn}+\lambda_{CE}{\mathcal{L}}_{CE}+\lambda_{SS}{\mathcal{L}}_{SS}+\lambda_{SI}{\mathcal{L}}_{SI}.$
@@
Based on the well-trained neural dialogue dynamics $f_{\theta},g_{\theta}$ and the NBF $\phi_{k}$ at each turn $k$, we introduce the filtering-based steering as a multi-turn jailbreak defense method. Given each state $x_{k-1}$ at each turn $k=1,2,\dots,K$, the NBF-based steering filters out all harmful attack queries $\hat{u}_{k}$ where $h(x_{k-1},\hat{u}_{k})+\eta\geq 0$ among jailbreaking context set ${\mathcal{U}}_{k-1}$ from attack methods, resulting in the safe query $u_{k}$ satisfying both Equation˜9 and $\phi_{k}(x_{k-1})=h(x_{k-1},u_{k})<0$. Therefore, given the well-trained neural dialogue dynamics and NBF, with unsafe queries being filtered out by the proposed NBF-based steering, invariant safety condition in Equation˜10 will be satisfied and LLM safety will not be jailbroken by Corollary 4.2.1. Empirically, the worst-case query candidates to be filtered come from jailbreaking attack methods (ren2024derail, ; russinovich2024great, ; li2024llm, ), which are the most harmful reasonable queries along the context set ${\mathcal{S}}_{context}$ at each turn in Definition 3.1. An example of Crescendo attack (russinovich2024great, ) can be found in Figure˜4, where $h(x_{k-1},u_{k})$ serves as a safety index to filter out the adversarial harmful query.
@@
LLMs
LLMs w/ steering
Attack Success Rate (ASR, $\downarrow$) Helpfulness ($\uparrow$) Over-Refusal ($\downarrow$)
[…]
GPT-3.5-turbo 0.585 0.560 0.785 8.00 0.078
GPT-3.5-turbo + steering 0.040 0.235 0.375 7.59 0.078
GPT-4o 0.600 0.565 0.725 9.35 0.004
GPT-4o + steering 0.035 0.260 0.325 8.77 0.026
o1 0.510 0.445 0.530 9.22 0.039
o1 + steering 0.090 0.280 0.210 8.83 0.057
Claude 3.5 Sonnet 0.200 0.215 0.095 9.14 0.052
Claude 3.5 Sonnet + steering 0.040 0.120 0.045 8.61 0.074
Table 1: Safety, helpfulness and over-refusal of closed-source LLMs before and after NBF steering.
@@ 5 Experiments
In this section, we aim to answer two questions: How does the proposed NBF-based safety steering perform as a defense method against different multi-turn LLM jailbreaking attacks on different LLMs? How is the proposed method influenced by steering threshold and training losses in terms of both safety and helpfulness? We answer the first question in Section˜5.2 and the second one in Section˜5.3, following the experimental setup. More details and results can be found in Appendix˜B.
@@ 5.1 Experimental Setup
To collect adversarial conversation data and safety labels for model training, we first generate diverse multi-turn jailbreaking attacks (ren2024derail, ; russinovich2024great, ; li2024llm, ) and responses of GPT-3.5-turbo based on training queries from Circuit Breakers (zou2024improving, ). Following ren2024derail , we adopt GPT-4o as the LLM safety judge (qi2023fine, ) to obtain safety scores ($1\sim 5$) as labels for each turn. Based on the collected multi-turn jailbreaking queries and responses, we first obtain the embeddings using the state-of-the-art pretrained embedding models all-mpnet-base-v2 (song2020mpnet, ; reimers2019sentence, ) (default) and all-distilroberta-v1 (sanh2019distilbert, ; nikolaev2023representation, ), and use the embeddings to train the neural dialogue dynamics $f_{\theta},g_{\theta}$ based on Equation˜4 for 200 epochs with Adam and learning rate $1e^{-4}$. We let the state dimension be 768, and $f_{\theta},g_{\theta}$ can be parameterized by 3-layer ReLU-based MLPs with the dimension shape of 1536-512-512-768. We then parameterize the safety predictor $h$ using 3-layer ReLU-based MLPs with the dimension shape of 1536-32-32-5 based on the pretrained neural dialogue dynamics $f_{\theta},g_{\theta}$. Given safety score labels, the predictor-based neural barrier function is learned based on Equations˜6, 11 and 12 for 200 epochs with Adam and learning rate $1e^{-3}$. The steering threshold is $\eta=0$ and the number of non-invariant turns $\kappa$ is 3 by default during model training.
Attack Success Rate (ASR, $\downarrow$) original
+ system
@@
llama-3-8b-
instruct
ActorAttack 0.425 0.280 0.070 0.120 0.040
@@
Phi-4
ActorAttack 0.405 0.370 0.100 0.080 0.015
@@
Table 2: Multi-turn safety comparison with defense baselines, highlighting the best and the runner-up.
Helpfulness and Over-Refusal Rate original
+ system
prompt
+ LoRA
SFT
+ NBF steering
($\eta=0$)
+ NBF steering
($\eta=1e^{-3}$)
llama-3-8b-
instruct
MMLU ($\uparrow$) 66.00 65.66 63.34 64.52 46.65
MTBench ($\uparrow$) 7.96 8.13 7.52 7.90 7.42
XSTest ($\downarrow$) 0.078 0.178 0.217 0.087 0.096
JailbreakBench-Benign ($\downarrow$) 0.34 0.49 0.34 0.4 0.4
Phi-4
MMLU ($\uparrow$) 78.49 78.67 76.77 76.68 56.09
MTBench ($\uparrow$) 8.23 8.59 8.06 8.18 7.76
XSTest ($\downarrow$) 0.087 0.052 0.139 0.087 0.100
JailbreakBench-Benign ($\downarrow$) 0.18 0.17 0.22 0.26 0.27
Table 3: Helpfulness and over-refusal comparison with baselines with the best and the runner-up.
Evaluation and baselines.
We apply NBF-based safety steering on different LLMs, including GPT-3.5 (gpt-3.5-turbo-0125) (openai2023gpt35, ), GPT-4o (gpt-4o-2024-08-06) (openai2024gpt4o, ), o1 (o1-2024-12-17) (openai2024o1, ), Claude-3.5 (claude-3-5-sonnet-20241022) (anthropic2024claude35, ), Llama-3-8b-instruct and Llama-3.1-80b (dubey2024llama, ), and Phi-4 (abdin2024phi, ). We evaluate the defense performance against state-of-the-art multiturn jailbreaking attacks (ActorAttack (ren2024derail, ), Crescendo (russinovich2024great, ), and Opposite-day (li2024llm, )) based on Harmbench dataset (mazeika2024harmbench, ; ren2024derail, ), using Attack Success Rate (ASR) metric as the ratio of successful jailbreaks judged by GPT-4o (qi2023fine, ) following ren2024derail . All temperatures are set to be 0.7. Besides, we evaluate the helpfulness using MMLU (hendrycks2020measuring, ) and MTBench (zheng2023judging, ), where MMLU assesses a model’s general knowledge across various subjects and reports the percentage of accuracy, while MTBench specifically evaluates an LLM’s ability to handle multi-turn conversations in dialogue scenarios with scores from 1 to 10. We further compare the over-refusal rate on XSTest and the benign behaviors of JailbreakBench, showing how conservative the proposed NBF-based steering will be under benign queries directly related to safety. We implement two defense baselines in open-source LLMs: supervised fine-tuning (SFT) with LoRA (hu2021lora, ; zheng2024llamafactory, ) and prompt-based steering from Llama-2-Chat (llama2prompt, ; touvron2023llama, ). Additionally, for a fair comparison with other filtering-based LLM guardrails, we compare ours with the lightweight baseline guardrails of OpenAI Moderation (markov2023holistic, ), ShieldGemma (zeng2024shieldgemma, ) and LLaMA Guard (inan2023llama, ), and report the F1 score of the prompt harmfulness detection (liu2025guardreasoner, ) under HarmBench (mazeika2024harmbench, ), AegisSafetyTest (ghosh2024aegis, ) and WildGuardTest (han2024wildguard, ).
F1 of prompt harmfulness detection ($\uparrow$) Model Size HarmBench AegisSafetyTest WildGuardTest
OpenAI Moderation Unknown 0.096 0.319 0.121
ShieldGemma 2B 0.118 0.075 0.094
LLaMA Guard 7B 0.672 0.741 0.560
MPNet-based NBF (Ours) 115M 0.811 0.748 0.572
DistilRoBERTa-based NBF (Ours) 87M 0.848 0.740 0.617
Table 4: F1 score comparison of current guardrails and ours regarding prompt harmfulness detection.
@@ 5.2 Result Comparison regarding Safety and Helpfulness on Multiple LLMs
Multi-turn attack comparison.
From Table˜1, we can see that our proposed NBF-based steering with threshold $\eta=1e^{-3}$ can significantly reduce ASR against all multi-turn jailbreaking attacks on different LLMs, showing the effectiveness and strong generalizability of the proposed neural dialogue dynamics and barrier function to unseen LLMs. Compared to the defense baselines of prompt-based steering and LoRA SFT in Table˜2, our steering defense has the lowest ASR under the threshold $\eta=1e^{-3}$. Notably, LoRA SFT can even lead to a higher ASR due to the presence of benign data in the fine-tuning process, which is consistent with the findings of qi2023fine ; he2024s ; qi2025safety .
Generalizability over unseen multi-turn attacks.
To evaluate the generalizability of our steering method, we further train the safety predictor with fewer attack methods, and test and compare ASR of these unseen attack methods and MTBench score with the SFT baseline. From the tables below, we can see that even though the NBF is trained without ActorAttack and Opposite-day attack data, our safety predictor can generalize better to these unseen attacks and yield a better trade-off of utility compared to the SFT method.
Attack success rate
and helpfulness
Llama3-8b-instruct Phi-4
original + SFT + steering original + SFT + steering
ActorAttack 0.425 0.180 0.175 0.405 0.160 0.155
Opposite-day 0.405 0.210 0.150 0.330 0.210 0.150
MTBench 7.96 7.54 7.80 8.23 7.92 7.93
Table 5: Generalizability of models trained without ActorAttack and Opposite-day.
Helpfulness and over-refusal comparison.
We compare the helpfulness and over-refusal in Tables˜1 and 3 to show the trade-off caused by safety steering. In Table˜3, it can be seen that the prompt-based steering can contribute to helpfulness while other defense methods tend to slightly compromise helpfulness. Specifically, our steering with threshold $\eta=0$ is more helpful than LoRA SFT. Although a stronger steering with threshold $\eta=1e^{-3}$ will cause a larger drop in MMLU due to filtering out unseen general knowledge without context, it still maintains satisfactory multi-turn conversation ability with low over-refusal rate over MTBench and XSTest in Tables˜1 and 3.
Comparison of prompt harmfulness detection.
From Table˜4, it can be seen that the proposed NBF-based safety filtering outperforms the baselines in all three benchmarks. Specifically, F1 scores of both MPNet-based and DistilRoBERTa-based NBF filtering are significantly higher than the others on HarmBench. Besides, our model sizes including different pretrained embedding models are significantly smaller than LLM guard baselines, validating the effectiveness of the proposed method as lightweight add-on post-training guardrails.
@@ 5.3 Ablation Study
Since the size of safe set ${\mathcal{X}}_{k}$ is controlled by the steering threshold $\eta>0$ in Equation˜7, the larger $\eta$ is, the stronger safety steering will be. From Figure˜5, we can see the Perato front induced by the steering threshold, showing the trade-off between helpfulness and safety. With the additional safe system prompt, LLMs can be safer and more helpful, but the attack success rate is still high. Compared to LoRA SFT, the proposed safety steering has a better trade-off and flexibility, being either more helpful given safety or safer given helpfulness.
@@
We compare the steering performance of safety and helpfulness based on NBF trained without either safe set loss ${\mathcal{L}}_{SS}$ of Equation˜11 or safety invariance loss ${\mathcal{L}}_{SI}$ Equation˜12 in Table˜6 under safety steering threshold $\eta=0$. We can find that with respect to safety, ablating ${\mathcal{L}}_{SS}$ or ${\mathcal{L}}_{SI}$ will mostly increase ASR, showing that these two proposed losses are essential to train neural barrier functions. As a trade-off, adding ${\mathcal{L}}_{SS}$ or ${\mathcal{L}}_{SI}$ will slightly hurt single-turn MMLU helpfulness, while multi-turn helpfulness on MTBench will be better with safety invariance loss ${\mathcal{L}}_{SI}$.
LLMs ${\mathcal{L}}_{SS}$ ${\mathcal{L}}_{SI}$ Embedding Attack Success Rate (ASR, $\downarrow$) Helpfulness ($\uparrow$) Over-refusal Rate($\downarrow$)
ActorAttack Crescendo Opposite-day MMLU MTBench XSTest JailbreakBench
GPT-3.5-
turbo
$\checkmark$ $\times$ MPNet 0.385 0.555 0.705 67.64 7.68 0.130 0.27
$\times$ $\checkmark$ MPNet 0.205 0.555 0.740 67.79 8.01 0.126 0.22
$\checkmark$ $\checkmark$ MPNet 0.135 0.430 0.655 66.24 7.93 0.078 0.30
$\checkmark$ $\checkmark$ DistilRoBERTa 0.240 0.375 0.505 57.08 7.46 0.148 0.24
Llama-3-
8b-instruct
$\checkmark$ $\times$ MPNet 0.245 0.385 0.370 65.80 7.66 0.165 0.38
$\times$ $\checkmark$ MPNet 0.175 0.470 0.310 65.96 7.97 0.109 0.32
$\checkmark$ $\checkmark$ MPNet 0.120 0.360 0.310 64.52 7.90 0.104 0.40
$\checkmark$ $\checkmark$ DistilRoBERTa 0.135 0.270 0.115 55.58 7.50 0.143 0.37
Table 6: Effectiveness of the proposed safe set loss ${\mathcal{L}}_{SS}$ in Equation˜11 and safe invariance loss ${\mathcal{L}}_{SI}$ in Equation˜12 with different pretrained embedding models under filtering threshold $\eta=0$.
Influence of different pretrained embeddings.
Since our neural dialogue dynamics and barrier function are based on pretrained language embedding models, we compare the results under different pretrained embeddings in Tables˜4 and 6. Although the model size of DistilRoBERTa-based NBF is smaller, it induces more aggressive safety boundary in the embedding space, resulting in lower ASR against multi-turn attacks and larger over-refusal rate on XSTest in Table˜6, and better prompt harmfulness detection performance in Table˜4. In contrast, MPNet-based NBF gives better trade-off between safety and helpfulness, which provides more flexibility with different filtering thresholds.